OpenStudy (anonymous):

A function value and a quadrant are given. Find the other 5 function values. Give exact answers. cot theta=2, Quadrant 1. Okay so once I did the math, for sin, I got squareroot5/5 Cos theta=2squareroot5/5 and tan as 1/2. It keeps saying my answer is wrong. What am I doing that's incorrect? If it's in quadrant 1, the values should all be positive right?

OpenStudy (aum):



OpenStudy (aum):

Does it say anywhere how they want the answer: fraction, decimal, etc.

OpenStudy (anonymous):

It says to use radicals as needed. So fractions would be the preferred method.

OpenStudy (aum):

You have the correct answers. Try without rationalizing the denominator. \(\large \sin(\theta) = \frac{1}{\sqrt{5}}\)

OpenStudy (anonymous):

On the program I use, we're not allowed to keep the radicals on the denominator. So I'm not sure what's wrong with my math.

OpenStudy (aum):

You have the correct answers. Maybe you are entering it wrong. Are you using the word square root instead of clicking on something that will insert the radical sign?

OpenStudy (aum):

Is it saying even tan(theta) = 1/2 is wrong?

OpenStudy (anonymous):

WOW. Okay so I just checked it and what I inputted and what the program listed as the correct answers are the same. Yet it marked me as wrong. I think I'm just going to show this to my professor and ask if she can give me some form of credit. Thank you for checking my work!

OpenStudy (aum):

You are welcome.